Height=h Base=h+7 Hypot=13 Find height and length
Answer by edjones(8007)

h^2+(h+7)^2=13^2
a^2+b^2=c^2
h^2+h^2+14h+49=169
2h^2+14h-120=0
2(h^2+7h-60)=0
(h+12)(h-5)=0
h=-12 (discard) or h=5
height=5
base=5+7=12
Ed